Harold Reese must choose between two bonds: Bond X pays $62 annual interest and has a market value of $915. It has 14 years to maturity. Bond Z pays $52 annual interest and has a market value of $900. It has six years to maturity. Assume the par value of the bonds is $1,000. a. Compute the current yield on both bonds. (Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places.)
Answer #1

The current yield to maturity =Annual coupon/Market price

Bond X =62/915 =6.78%

Bond Z =52/900 =5.78%